OpenStudy (thatcooolswag):

Rewrite the following expression without grouping Symbols: 7(x + 9)

OpenStudy (greatlife44):

@thatcooolswag have you heard of the distributative property?

OpenStudy (thatcooolswag):

yah

OpenStudy (greatlife44):

\[7(x+9) = ? \] see those ( ) we must multiply whatever number outside the parenthesis by everything inside it

OpenStudy (greatlife44):

so show me what you think it might be?

OpenStudy (greatlife44):

@thatcooolswag

OpenStudy (thatcooolswag):

my teacher never really taught us for this lesson

OpenStudy (thatcooolswag):

@greatlife44

OpenStudy (greatlife44):

\[7(x+9) = 7*x+7*9\]

OpenStudy (greatlife44):

@thatcooolswag Let me give you another problem to do . Please solve this for me 4(x+2)

OpenStudy (thatcooolswag):

my typing is messed up

OpenStudy (greatlife44):

@thatcooolswag what's the answer?

OpenStudy (greatlife44):

Just try :)